What is the potential difference between two points in a circuit if 200J of electrical energy are changed to other forms of energy when 25 coulombs of electric charge pass? If the charge flows in 10 seconds, what is the current?
1
Expert's answer
2018-02-26T09:19:07-0500
The work done by electric field
W=qV
So, the potential difference
V=W/q=200/25=8 V
The current
I=q/t=25/10=2.5 A
Answers:
8 V
2.5 A